A pendulum takes 10 seconds to swing






through 2 complete cycles.







a. How long does it take to complete






one cycle?






b. What is its period?






c. What is its frequency?






d. What is the angular frequency?






e. What does the position N represents?

1
Expert's answer
2022-02-05T10:26:26-0500

a. How long does it take to complete one cycle? 5 s because each cycle has equal duration

b. What is its period? The period is 5 s because it's the time a cycle takes to complete.

c. What is its frequency? The frequency is the inverse of period, or 1/5=0.2 Hz.

d. What is the angular frequency? "\omega=2\pi f=2\pi·0.2=1.26\text{ rad/s}."

e. What does the position N represents? It represents the number of oscillations.